SubtoNOT:

A diesel engine's piston compresses 16 cm3 of fuel-air mixture into 1 cm3. The pressure changes from 1 atmosphere to 48 atmospheres. If the initial temperature of the gas was 305 K, what was the final temperature? Note: As long as the units for pressure and volume are the same on both sides of the equation, they will cancel. Temperature, however, must be in units of kelvin. Be sure to use the proper number of significant figures.

SubtoNOT:

jay do not

Jayisbetter:

To solve this problem, you can use the ideal gas law, which relates the initial and final conditions of a gas when no chemical reactions occur. The ideal gas law is given by: PV = nRT Where: P = Pressure (in atmospheres) V = Volume (in cm³) n = Number of moles R = Ideal gas constant (8.31 J/(mol·K)) T = Temperature (in kelvin) In your problem, we have an initial condition and a final condition, and we want to find the final temperature, T_final. To do this, we can use the ideal gas law for both initial and final conditions and set them equal to each other: (P_initial * V_initial) = (P_final * V_final) Here's what we know: P_initial = 1 atmosphere V_initial = 16 cm³ P_final = 48 atmospheres V_final = 1 cm³ T_initial = 305 K We need to find T_final. First, let's calculate the number of moles (n) using the initial condition: n = (P_initial * V_initial) / (R * T_initial) n = (1 atm * 16 cm³) / (8.31 J/(mol·K) * 305 K) n ≈ 0.00615 moles Now that we know the number of moles (n), we can use the ideal gas law to find T_final: (P_final * V_final) = n * R * T_final T_final = (P_final * V_final) / (n * R) T_final = (48 atm * 1 cm³) / (0.00615 moles * 8.31 J/(mol·K)) T_final ≈ 878.7 K So, the final temperature is approximately 878.7 kelvin. Be sure to keep the proper number of significant figures, which in this case is one decimal place.

SubtoNOT:

ai gets lots of things wrong (skull)

@jayisbetter wrote:
To solve this problem, you can use the ideal gas law, which relates the initial and final conditions of a gas when no chemical reactions occur. The ideal gas law is given by: PV = nRT Where: P = Pressure (in atmospheres) V = Volume (in cm³) n = Number of moles R = Ideal gas constant (8.31 J/(mol·K)) T = Temperature (in kelvin) In your problem, we have an initial condition and a final condition, and we want to find the final temperature, T_final. To do this, we can use the ideal gas law for both initial and final conditions and set them equal to each other: (P_initial * V_initial) = (P_final * V_final) Here's what we know: P_initial = 1 atmosphere V_initial = 16 cm³ P_final = 48 atmospheres V_final = 1 cm³ T_initial = 305 K We need to find T_final. First, let's calculate the number of moles (n) using the initial condition: n = (P_initial * V_initial) / (R * T_initial) n = (1 atm * 16 cm³) / (8.31 J/(mol·K) * 305 K) n ≈ 0.00615 moles Now that we know the number of moles (n), we can use the ideal gas law to find T_final: (P_final * V_final) = n * R * T_final T_final = (P_final * V_final) / (n * R) T_final = (48 atm * 1 cm³) / (0.00615 moles * 8.31 J/(mol·K)) T_final ≈ 878.7 K So, the final temperature is approximately 878.7 kelvin. Be sure to keep the proper number of significant figures, which in this case is one decimal place.

Fresno:

✔️

Rosee5656:

coolozs